# When Models Don’t Collapse: On the Consistency of Iterative MLE

### v1

@@ Abstract
The widespread use of generative models has created a feedback loop, in which each generation of models is trained on data partially produced by its predecessors. This process has raised concerns about model collapse: A critical degradation in performance caused by repeated training on synthetic data. However, different analyses in the literature have reached different conclusions as to the severity of model collapse. As such, it remains unclear how concerning this phenomenon is, and under which assumptions it can be avoided. To address this, we theoretically study model collapse for maximum likelihood estimation (MLE), in a natural setting where synthetic data is gradually added to the original data set. Under standard assumptions (similar to those long used for proving asymptotic consistency and normality of MLE), we establish non-asymptotic bounds showing that collapse can be avoided even as the fraction of real data vanishes. On the other hand, we prove that some assumptions (beyond MLE consistency) are indeed necessary: Without them, model collapse can occur arbitrarily quickly, even when the original data is still present in the training set. To the best of our knowledge, these are the first rigorous examples of iterative generative modeling with accumulating data that rapidly leads to model collapse.
@@ 1 Introduction
Nowadays, generative models are filling the internet with synthetic data. At the same time, web-scraped content remains a common source for training newer models. As a result, one can envision a feedback loop, in which each generation of models is trained in part on content produced by its predecessors. Such a loop may persist indefinitely, eventually causing a substantial portion of training data to be synthetic. When this happens, repeated training on artificial content can amplify existing model biases and degrade performance across generations. This phenomenon has recently been termed Model Collapse: A critical degradation in performance caused by repeated training on synthetic data (Shumailov et al., 2024; Bertrand et al., 2024; Dohmatob et al., 2024a; Gerstgrasser et al., 2025).
@@
To help clarify this picture, we theoretically study the behavior of maximum likelihood estimation (MLE) under iterative training with synthetic data. We focus on a natural and practically motivated setting (also studied by (Alemohammad et al., 2024; Gerstgrasser et al., 2025; Dey and Donoho, 2024)), where we initially have $n$ samples from some ground-truth distribution, and at each iteration $T$, the latest generation model generates $n$ new samples that are then accumulated with all previous data and used to train the next model.
@@
Recently, Dey and Donoho (2024) analyzed a similar setting for exponential families, and showed that for any fixed $T$ and in the limit as $n\to\infty$, the error of the generation-$T$ model is degraded by at most a universal multiplicative constant compared to the initial estimator trained solely on real data. However, since their guarantees are asymptotic in $n$ while $T$ is fixed, they do not quantify how the performance depends on the proportion of synthetic versus real data. This is because in their setting, there is always a constant fraction of training data that is real. Therefore, it is difficult to deduce from their results how concerning model collapse may be as the fraction of real data decreases.
@@
In contrast, in this work, we prove in Thm. 4.1 non-asymptotic guarantees that remain valid even as the fraction of real data approaches zero. Under standard regularity and smoothness assumptions (of the kind long used to prove asymptotic consistency and normality of MLE), we show that as long as the number of samples per iteration is at least polylogarithmic in the number of iterations, iterative MLE is consistent (meaning that it converges to the ground truth model as the sample size increases). These findings offer a sharper theoretical understanding of when model collapse can be avoided.
@@
We complement this result with negative ones, that illustrate what can go wrong when these assumptions are violated. In particular, we construct families of distributions for which MLE is consistent when trained on real data, but nonetheless suffers from collapse when synthetic data is iteratively accumulated. Our negative results come in two flavors. In Thm. 5.1, for any fixed sample size $n$, we construct a family of distributions such that the first iteration gives an excellent approximation to the real distribution, but even the second does not with constant probability. Next, in Thm. 5.2, we show that there exists a family of distributions such that for any $n$, model collapse will eventually occur, after a number of iterations which grows arbitrarily slowly with $n$.
@@
To the best of our knowledge, Thm. 5.1 and Thm. 5.2 are the first rigorous examples of iterative generative modeling with accumulating data that rapidly lead to model collapse. Recently, it has been suggested that model collapse does not occur when data accumulates across iterations (Gerstgrasser et al., 2025; Dey and Donoho, 2024; Schaeffer et al., 2025). Our results show that such claims can only be true under structural assumptions beyond MLE consistency.
@@ 2 Related Work
Among theoretical works, several settings have been studied that differ somewhat from the setting of this paper. A notable line of works considers linear regression, taking advantage of the closed-form expression for the least squares estimate (Dohmatob et al., 2024a; Gerstgrasser et al., 2025; Dohmatob et al., 2025). These works focus on discriminative models, where previous models are used to label new data, not synthetically generate new data as in this paper.
@@ 3.1 Notation
[…]
@@ 3.2 Iterative Maximum Likelihood Estimation
Given a dataset $X\subseteq\mathcal{X}$ the MLE trained on $X$ is given by
@@
In the above definition, it is not immediately clear why the MLE exists or if it is unique. Existence is known to hold under mild assumptions, and throughout the proofs, we will explicitly show existence whenever necessary. Regarding uniqueness, the MLE may not be unique in general. However, under mild assumptions, it is known that the MLE converges to the real parameters $\theta^{\star}$ (e.g. (Wald, 1949)), and that given sufficiently many samples, the log-likelihood is strictly concave in a neighborhood of $\theta^{\star}$. As such, asymptotically, the MLE is expected to be unique. Nevertheless, formally treating this typically introduces unnecessary and undesired complications to the analysis. It is therefore standard to simply assume that the MLE is unique whenever it exists (e.g. (Lehmann and Casella, 2006)). We follow (Bertrand et al., 2024) in making a similar, but slightly milder assumption that if there are multiple parameter vectors maximizing the log-likelihood, the argmax may choose the one that is closest to a given reference point. This is, of course, made explicit in the proofs.
@@
Assumption  1. B is necessary to quantify the distance between distributions $p_{\theta}$, $p_{\theta^{\prime}}$ using $\left\|\theta-\theta^{\prime}\right\|$. Note that one can always satisfy Assumption  1. B by removing duplicates from $P_{\Theta}$, or by considering the quotient topology as in Redner (1981). Assumption  1. C avoids pathologies and ensures that $\log p_{\theta}(x)$ is well-defined throughout the iterative sampling process. In distributions modeled using neural networks, probabilities are often given by applying a softmax, ensuring that they are always positive and thus satisfying Assumption  1. C.
@@ 4 Consistency of Iterative MLE
Under the same assumptions as in Thm. 4.1, convergence of parameters also implies convergence in KL-Divergence and convergence in total variation (TV) distance. We refer the reader to Appendix A.1 for background and details. In particular, for a suitable absolute constant $C>0$, Thm. 4.1 implies
@@
We now detail some ways in which Thm. 4.1 differs from past results on model (non)-collapse. In Bertrand et al. (2024), synthetic data does not accumulate across iterations, and for each iteration $t\in[T]$, most of the training data used to train $\theta^{(t)}$ is real. The maximal fraction of synthetic data was increased in the follow-up work Ferbach et al. (2024) when assuming access to the full distribution (i.e. $n=\infty$). Similarly, Dey and Donoho (2024) first fix $T$ and then analyze the limit of $n\to\infty$. They do not provide finite sample guarantees that quantify the dependence between $T$ and $n$. Seddik et al. (2024) bounded the expected value of the TV distance for distributions over finite vocabularies. When the amount of synthetic data is sufficiently large relative to the vocabulary size, their bound scales as $\mathcal{O}\left(\sqrt{k}/n\right)$ where $k$ is the total amount of synthetic data. In the data accumulation setting, $k=(T-1)n$, in which case the bound becomes $\mathcal{O}\left(\sqrt{T/n}\right)$.
4.1 Proof Sketch of Thm. 4.1
We provide here the proof intuition for Thm. 4.1, and refer the reader to Appendix D for the rigorous proof.
@@
We note that Thm. 4.1 cannot be obtained naively as a direct consequence of the Proposition C.1. Extending Proposition C.1 to a bound on $\left\|\theta^{(T)}-\theta^{(0)}\right\|$ using the triangle inequality leads to a suboptimal dependence on $T$, since it doesn’t take into account cancellations from iteration to iteration. Instead, as we will show in the following paragraph, Proposition C.1 will be used to ensure that for large $n$, $\theta^{(t+1)}$ will be sufficiently close to $\theta^{(t)}$ to enable Taylor expanding the log likelihood around it. This idea draws inspiration from the asymptotic normality analysis of MLE (Cramér, 1946; Lehmann and Casella, 2006).
@@
The expected value of $\nabla\ell_{t}\left(\theta^{(t)}\right)$ (conditioned on $\theta^{(t)}$ can be shown to be zero, so that the first term forms a martingale, which allows us to bound the norm essentially as if all samples were independent. Since each $\nabla\ell_{t}$ is scaled by $\frac{1}{t+1}$, the variance scales as $\frac{1}{(t+1)^{2}}$. So the variance of the sum can be upper bounded as $\sum_{t=1}^{T}\frac{1}{t^{2}}\leq\sum_{t=1}^{\infty}\frac{1}{t^{2}}\leq\frac{\pi^{2}}{6}$. In summary, we show
@@
The second term in Eq. (4.1) has to be treated differently, as correlations between $H_{t-1}$ and $\nabla\ell_{t}$ imply that each term is not necessarily mean-zero, and so the sum should be expected to have some dependence on $T$. This term somewhat complicates the proof, as bounding it requires knowing that $\left\|\theta^{(t)}-\theta^{({0})}\right\|$ is sufficiently small for all $t<T$. The proof thus works inductively, bounding this term from iteration to iteration. Roughly speaking, in the end, we show that
@@ 5 Necessity of Structural Assumptions
Thm. 4.1 provides conditions under which the iterative MLE retains good performance, even if the proportion of synthetic data approaches $1$. Clearly, this cannot always be true. In particular, there are well-known examples of families of distributions on which even standard MLE is inconsistent: Namely it will not converge to the ground-truth parameters as the sample size increases, even when trained purely on real data (e.g. (Bahadur, 1958; Ferguson, 1982; Le Cam, 1990)). In such situations, the whole question of model collapse is rather meaningless. Thus, a natural (informal) follow-up question is the following: In the setting where synthetic data is added to the real dataset in each iteration, is there a family of distributions that is sufficiently well-behaved for MLE to be asymptotically consistent (when trained on real data), but still exhibits rapid model collapse? In other words, do there exist cases where the MLE can learn the real distribution, and yet model collapse still occurs when applying MLE iteratively?
In this section, we show that the answer is yes, and demonstrate different settings in which model collapse can occur when the conditions of Thm. 4.1 are not satisfied. To the best of our knowledge, these are the first rigorous examples of iterative generative modeling with accumulating data that rapidly leads to model collapse.
We emphasize that, following the rest of the paper, we focus here on a setting where synthetic data iteratively accumulates on top of the real data. A different model collapse setting studied in some previous works is when at each iteration, MLE is performed purely on synthetic data generated by the latest model. In such a setting, the real training data disappears already after a single iteration, and it has been shown to lead to model collapse even for very well-behaved distributions such as Gaussians (Shumailov et al., 2024). It has recently been suggested that if data is added rather than replaced (as in our setting), the extent to which iterative MLE performance degrades is limited (Gerstgrasser et al., 2025; Dey and Donoho, 2024; Schaeffer et al., 2025). We show here that this can be true only if further assumptions are made, beyond just MLE consistency (as we do in Thm. 4.1).
@@
Note that we use here convergence in total variation, rather than convergence in parameters as in Thm. 4.1. The reason is that to establish our negative results, we have to make use of distributions that do not follow the assumptions of Thm. 4.1, and in particular do not satisfy the smoothness assumptions there. Without smoothness, parametric convergence and convergence of distributions are no longer equivalent in general. Thus, using a probability metric such as total variation is more natural in our setting, as we are ultimately interested in approximating the ground-truth distribution.

### v2

@@ Abstract
The widespread use of generative models has created a feedback loop, in which each version of a model is trained on data partially produced by its predecessors. This process has raised concerns about model collapse: A critical degradation in performance caused by repeated training on synthetic data. However, different analyses in the literature have reached different conclusions as to the severity of model collapse. As such, it remains unclear how concerning this phenomenon is, and under which assumptions it can be avoided. To address this, we theoretically study model collapse for maximum likelihood estimation (MLE), in a natural setting where synthetic data is gradually added to the original data set. Under standard assumptions (similar to those long used for proving asymptotic consistency and normality of MLE), we establish non-asymptotic bounds showing that collapse can be avoided even as the fraction of real data vanishes. On the other hand, we prove that some assumptions (beyond MLE consistency) are indeed necessary: Without them, model collapse can occur arbitrarily quickly, even when the original data is still present in the training set. To the best of our knowledge, these are the first rigorous examples of iterative generative modeling with accumulating data that rapidly leads to model collapse.
@@ 1 Introduction
Generative models such as large language models (LLMs) and diffusion models are increasingly filling the internet with synthetic data. At the same time, web-scraped content remains a common source for training newer models. This creates a feedback loop in which each version of a model is trained partly on outputs of all past models, and synthetic data gradually dominates future training sets. Because identifying and filtering artificially generated content at scale is not always feasible, existing biases and artifacts in the models can become increasingly embedded in the training data and amplified from model to model. This phenomenon has recently been termed Model Collapse: A critical degradation in performance caused by repeated training on synthetic data (Shumailov et al., 2024; Bertrand et al., 2024; Dohmatob et al., 2024a; Gerstgrasser et al., 2025).
@@
To help clarify this picture, we theoretically study the behavior of maximum likelihood estimation (MLE) under iterative training with synthetic data. We focus on a natural and practically motivated setting (also studied by Alemohammad et al. (2024); Gerstgrasser et al. (2025); Dey and Donoho (2024)), where we initially have $n$ samples from some ground-truth distribution, and at each iteration $T$, the latest model generates $n$ new samples that are then accumulated with all previous data and used to train the next model.
@@
Recently, Dey and Donoho (2024) analyzed a similar setting specifically for distributions arising from exponential families, and showed that for any fixed $T$ and in the limit as $n\to\infty$, the error of the iteration-$T$ model is degraded by at most a universal multiplicative constant compared to the initial estimator trained solely on real data. However, their guarantees are asymptotic in $n$ while $T$ is fixed. Therefore, they do not quantify how the performance depends on the proportion of synthetic versus real data, since the fraction of training data that is real is a constant (given by $1/T$) while $n$ grows to infinity. Therefore, it is difficult to deduce from their results how concerning model collapse may be as the fraction of real data decreases.
@@
In contrast, in this work, we prove in Theorem 4.1 non-asymptotic guarantees that remain valid even as the fraction of real data approaches zero. Under standard regularity and smoothness assumptions (of the kind long used to prove asymptotic consistency and normality of MLE), we show that as long as the number of samples per iteration is at least polylogarithmic in the number of iterations, iterative MLE is consistent (meaning that it converges to the ground truth model as the sample size increases). These findings offer a sharper theoretical understanding of when model collapse can be avoided.
@@
We complement this result with negative ones, that illustrate what can go wrong when these assumptions are violated. In particular, we construct families of distributions for which MLE is consistent when trained on real data, but nonetheless suffers from collapse when synthetic data is iteratively accumulated. Our negative results come in two flavors. In Theorem 5.1, for any fixed sample size $n$, we construct a family of distributions such that the first iteration gives an excellent approximation to the real distribution, but even the second does not with constant probability. Next, in Theorem 5.2, we show that there exists a family of distributions such that for any $n$, model collapse will eventually occur, after a number of iterations which grows arbitrarily slowly with $n$.
@@
To the best of our knowledge, Theorem 5.1 and Theorem 5.2 are the first rigorous examples of iterative generative modeling with accumulating data that rapidly lead to model collapse. Recently, it has been suggested that model collapse does not occur when data accumulates across iterations (Gerstgrasser et al., 2025; Dey and Donoho, 2024; Schaeffer et al., 2025). Our results show that such claims can only be true under structural assumptions beyond MLE consistency.
@@ 2 Related Work
Among theoretical works, several settings have been studied that differ somewhat from the setting of this paper. A notable line of works considers linear regression, taking advantage of the closed-form expression for the least squares estimate (Dohmatob et al., 2024a; Gerstgrasser et al., 2025; Dohmatob et al., 2025). These works focus on discriminative models, where previous models are used to label new data, not synthetically generate new data as in this paper. Moreover, the setting of (Dohmatob et al., 2025) is non-iterative, and they analyze the test error when the training data contains some samples that are labeled from a linear predictor drawn from a “bad” synthetic distribution that differs from the real one. This is a key difference from works such as Dey and Donoho (2024) that analyzed a setting where data gradually accumulates.
@@ 3.1 Notation
We use bold-faced font to denote vectors, e.g. $\mathbf{x}\in\mathbb{R}^{d}$, and denote by $\left\|\mathbf{x}\right\|$ the Euclidean norm. We let $n]:=\{1,\dots,n\}$. Unless otherwise stated, $\left\|\cdot\right\|$ denotes the operator norm for matrices and 33rd-order tensors, where the latter is defined for a 33rd-order tensor $A$ as $\left\|A\right\|=\sup_{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\neq\boldsymbol{0}}\frac{A(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3})}{\left\|\mathbf{v}_{1}\right\|\left\|\mathbf{v}_{2}\right\|\left\|\mathbf{v}_{3}\right\|}$. We use the standard big-O notation, with $\mathcal{O}(\cdot)$, $\Theta(\cdot)$ and $\Omega(\cdot)$ hiding absolute constants that do not depend on problem parameters. To specify constants that depend only on certain quantities, we may put these quantities in parentheses. For example, $C(K_{1},K_{2})$ would denote a constant that depends only on $K_{1},K_{2}$. For a given vector $\mathbf{v}$ and radius $r>0$, we let $B_{r}(\mathbf{v}):=\{\mathbf{u}:\left\|\mathbf{u}-\mathbf{v}\right\|\leq r\}$ be the closed ball of radius $r$ centered at $\mathbf{v}$. For a function $f(\theta)$, we write $\nabla^{2}f(\theta)$ for its Hessian, and $\nabla^{3}f(\theta)$ for its 33rd-order derivative tensor, meaning $\nabla^{3}f(\theta)]_{i,j,k}=\frac{\partial^{3}}{\partial\theta_{i}\partial\theta_{j}\partial\theta_{k}}f(\theta)$. For a matrix $A$, we denote by $\lambda_{\min}(A)$, $\lambda_{\max}(A)$ its minimal and maximal eigenvalues respectively.
@@ 3.2 Iterative Maximum Likelihood Estimation
Let $X$ be a dataset with elements belonging to $\mathcal{X}$, the MLE trained on $X$ is given by
@@
In the above definition, it is not immediately clear why the MLE exists or if it is unique. Existence is known to hold under mild assumptions, and throughout the proofs, we will explicitly show existence whenever necessary. Regarding uniqueness, the MLE may not be unique in general. However, under mild assumptions, it is known that the MLE converges to the real parameters $\theta^{\star}$ (e.g. (Wald, 1949)), and that given sufficiently many samples, the log-likelihood is strictly concave in a neighborhood of $\theta^{\star}$. As such, asymptotically, the MLE is expected to be unique. Nevertheless, formally treating this typically introduces unnecessary and undesired complications to the analysis. It is therefore standard to simply assume that the MLE is unique whenever it exists (e.g. (Lehmann and Casella, 2006)). We follow Bertrand et al. (2024) in making a similar, but slightly milder assumption that if there are multiple parameter vectors maximizing the log-likelihood, the argmax may choose the one that is closest to a given reference point. This is, of course, made explicit in the proofs.
@@
Assumption  1. B is necessary to quantify the distance between distributions $p_{\theta}$, $p_{\theta^{\prime}}$ using $\left\|\theta-\theta^{\prime}\right\|$. Note that one can always satisfy Assumption  1. B by removing duplicates from $P_{\Theta}$, or by considering the quotient topology as in Redner (1981). Assumption  1. C avoids pathologies and ensures that $\log p_{\theta}(x)$ is well-defined throughout the iterative sampling process. In distributions modeled using neural networks, probabilities are usually given by applying a softmax, ensuring that they are always positive and thus satisfying Assumption  1. C.
@@
We end by noting that the assumptions above are mostly satisfied (at least approximately) by neural networks. For example, the constant support assumption (Assumption  1. C) is trivially satisfied in standard architectures, since the softmax function widely used to assign probabilities is always non-zero. The smoothness assumptions can be satisfied in various settings, especially when using techniques such as weight decay, which are standard in modern LLM training. Of course, the exact bounds would depend on the architecture and the setting. In general, we believe that weakening these assumptions is quite feasible and is an interesting direction for future work.
@@ 4 Consistency of Iterative MLE
Under the same assumptions as in Theorem 4.1, convergence of parameters also implies convergence in KL-Divergence and convergence in total variation (TV) distance. We refer the reader to Appendix A.1 for background and details. In particular, for a suitable absolute constant $C>0$, Theorem 4.1 implies
@@
We now detail some ways in which Theorem 4.1 differs from past results on model (non)-collapse. In Bertrand et al. (2024), synthetic data does not accumulate across iterations, and for each iteration $t\in[T]$, most of the training data used to train $\theta^{(t)}$ is real. The maximal fraction of synthetic data was increased in the follow-up work Ferbach et al. (2024) when assuming access to the full distribution (i.e. $n=\infty$). Similarly, Dey and Donoho (2024) first fix $T$ and then analyze the limit of $n\to\infty$. They do not provide finite sample guarantees that quantify the dependence between $T$ and $n$. Seddik et al. (2024) bounded the expected value of the TV distance for distributions over finite vocabularies. When the amount of synthetic data is sufficiently large relative to the vocabulary size, their bound scales as $\mathcal{O}\left(\sqrt{k}/n\right)$ where $k$ is the total amount of synthetic data. In the data accumulation setting, k=(T1)nk=(T-1)n, in which case the bound becomes $\mathcal{O}\left(\sqrt{T/n}\right)$.
We note that while Theorem 4.1 considers a setting where the exact MLE is computable, we believe the theorem can naturally be extended to also accommodate an optimization error, where only an approximate MLE is available. Indeed, we empirically observe in Appendix G that for families of distributions for which an exact formula for the MLE is known, the results are robust to mild optimization error.
4.1 Proof Sketch of Theorem 4.1
We provide here the proof intuition for Theorem 4.1, and refer the reader to Appendix D for the rigorous proof.
@@
We note that Theorem 4.1 cannot be obtained naively as a direct consequence of the Proposition C.1. Extending Proposition C.1 to a bound on $\left\|\theta^{(T)}-\theta^{(0)}\right\|$ using the triangle inequality leads to a suboptimal dependence on $T$, since it doesn’t take into account cancellations from iteration to iteration. Instead, as we will show in the following paragraph, Proposition C.1 will be used to ensure that for large $n$, $\theta^{(t+1)}$ will be sufficiently close to $\theta^{(t)}$ to enable Taylor expanding the log likelihood around it. This idea draws inspiration from the asymptotic normality analysis of MLE (Cramér, 1946; Lehmann and Casella, 2006).
@@
The expected value of $\nabla\ell_{t}\left(\theta^{(t)}\right)$ (conditioned on $\theta^{(t)}$) can be shown to be zero, so that the first term forms a martingale, which allows us to bound the norm essentially as if all samples were independent. Since each $\nabla\ell_{t}$ is scaled by $\frac{1}{t+1}$, the variance scales as $\frac{1}{(t+1)^{2}}$. So the variance of the sum can be upper bounded as $\sum_{t=1}^{T}\frac{1}{t^{2}}\leq\sum_{t=1}^{\infty}\frac{1}{t^{2}}\leq\frac{\pi^{2}}{6}$. In summary, we show that with high probability
@@
The second term in Eq. (4.1) has to be treated differently, as correlations between $H_{t-1}$ and $\nabla\ell_{t}$ imply that each term is not necessarily mean-zero, and so the sum should be expected to have some dependence on $T$. This term somewhat complicates the proof, as bounding it requires knowing that $\left\|\theta^{(t)}-\theta^{({0})}\right\|$ is sufficiently small for all $t<T$. The proof thus works inductively, bounding this term from iteration to iteration. Roughly speaking, in the end, we show that with high probability
@@ 5 Necessity of Structural Assumptions
Theorem 4.1 provides conditions under which the iterative MLE retains good performance, even if the proportion of synthetic data approaches $1$. Clearly, this cannot always be true. In particular, there are well-known examples of families of distributions on which even standard MLE is inconsistent: Namely it will not converge to the ground-truth parameters as the sample size increases, even when trained purely on real data (e.g. (Bahadur, 1958; Ferguson, 1982; Le Cam, 1990)). In such situations, the whole question of model collapse is rather meaningless. Thus, a natural (informal) follow-up question is the following: In the setting where synthetic data is added to the real dataset in each iteration, is there a family of distributions that is sufficiently well-behaved for MLE to be asymptotically consistent (when trained on real data), but still exhibits rapid model collapse? In other words, do there exist cases where the MLE can learn the real distribution, and yet model collapse still occurs when applying MLE iteratively?
In this section, we show that the answer is yes, and demonstrate different settings in which model collapse can occur when the conditions of Theorem 4.1 are not satisfied. To the best of our knowledge, these are the first rigorous examples of iterative generative modeling with accumulating data that rapidly leads to model collapse.
We emphasize that, following the rest of the paper, we focus here on a setting where synthetic data iteratively accumulates on top of the real data. A different model collapse setting studied in some previous works is when at each iteration, MLE is performed purely on synthetic data generated by the latest model. In such a setting, the real training data disappears already after a single iteration, and it has been shown to lead to model collapse even for very well-behaved distributions such as Gaussians (Shumailov et al., 2024). It has recently been suggested that if data is added rather than replaced (as in our setting), the extent to which iterative MLE performance degrades is limited (Gerstgrasser et al., 2025; Dey and Donoho, 2024; Schaeffer et al., 2025). We show here that this can be true only if further assumptions are made, beyond just MLE consistency (as we do in Theorem 4.1).
@@
Note that we use here convergence in total variation, rather than convergence in parameters as in Theorem 4.1. The reason is that to establish our negative results, we have to make use of distributions that do not follow the assumptions of Theorem 4.1, and in particular do not satisfy the smoothness assumptions there. Without smoothness, parametric convergence and convergence of distributions are no longer equivalent in general. Thus, using a probability metric such as total variation is more natural in our setting, as we are ultimately interested in approximating the ground-truth distribution.